Suppose that the speed v (in ft/s) of a skydiver t seconds after leaping from a plane is given by the equation 𝑣 = 190(1 βˆ’ 𝑒 βˆ’ 0.168𝑑).

(a) Graph 𝑣 versus 𝑑.

(b) By evaluating an appropriate limit, show that the graph of 𝑣 versus 𝑑 has a horizontal asymptote 𝑣 = 𝑐 for an appropriate constant 𝑐.

(c) What is the physical significance of the constant 𝑐 in part (b)?


1
Expert's answer
2021-10-04T16:15:05-0400
"v(t)=190(1-e^{-0.168t})"

(a)








(b)


"\lim\limits_{t\to\infin}v(t)=\lim\limits_{t\to\infin}(190(1-e^{-0.168t}))=190(ft/s)"

The horizontal asymptote is "v=190\ ft/s."


(c) It is the terminal velocity of skydiver.